# Small Open Models Achieve Near-Parity with Large Models in Low-Resource Literary Translation at a Fraction of the Cost

### v1

@@ 1 Introduction
Translating literature into low-resource languages poses unique challenges, given both the scarcity of parallel data and the need to faithfully preserve stylistic nuance and cultural context. Romanian—a Latin language spoken by over 24 million people—remains notably underserved in the realm of high-quality literary translation resources. Existing machine translation (MT) benchmarks, such as WMT [^1] WMT stands for the Conference on Machine Translation (formerly “Workshop on Machine Translation”). The WMT organizes annual shared tasks for benchmarking machine translation systems, with its “news translation” task being the most widely used standard evaluation for translation of contemporary news articles. See: https://www.statmt.org/wmt23/., predominantly focus on informational or news text, thus providing little guidance for creative, narrative-driven translation tasks. The rapid advancement of large language models (LLMs) has, however, enabled synthetic text generation at unprecedented scale, opening up new opportunities for augmenting MT datasets [14, 19]. Notably, projects like TinyStories [4] have shown that even compact models can produce coherent narratives when trained on thoughtfully curated synthetic corpora. While these advances have transformed data generation and model pretraining, much of the prior work has centered on monolingual story generation in English [6, 4] or improvements to translation via back-translation [22], rather than the systematic creation of large-scale, bilingual literary datasets. As a result, there remain open questions about how both proprietary and open-source LLMs can be harnessed to create and evaluate high-quality literary translations in genuinely low-resource settings.
@@
Significant progress has recently been made toward building massive bilingual and multilingual resources—for example, CLIRMatrix [26] compiles parallel data across 139 languages for cross-lingual retrieval, and mT5 [33] pretrains a text-to-text transformer on a vast array of language pairs and domains. However, such resources are primarily oriented toward general-domain or retrieval-centric MT, with limited representation of literary content. Even in language pairs like English–Romanian, coverage is often restricted to news, web, or miscellaneous domains. To date, there is no open, large-scale, high-quality English–Romanian parallel corpus specifically curated for creative narratives or fables.
@@
TinyFabulist Translation Framework (TF2) addresses this gap. DS-TF2-EN-RO-3M is a 3-million-example synthetic dataset of English–Romanian parallel fables, accompanied by fine-tuned open-source translation models and a literary translation benchmark. It extends our earlier TinyFabulist-TF1 corpus of English-only fables [18] into a bilingual setting, explicitly targeting creative narrative translation and cultural adaptation. By combining parameter-efficient adaptation with narrative-aware evaluation, TF2 provides both a large-scale resource and a practical benchmark for studying literary MT in a genuinely low-resource language pair.
@@
While recent multilingual models such as NLLB-200 [27] and EuroLLM [16] have demonstrated impressive performance, these systems are generally optimized for breadth and general-domain translation. Their large parameter counts also limit practical deployment in cost-constrained or on-device scenarios. Moreover, our experiments show that such models, despite their coverage, often fall short in producing the nuanced style, coherence, and cultural adaptation required for professional literary translation (see Section 4).
Despite recent advances in both synthetic data creation and multilingual modeling, there remains a critical lack of resources and benchmarks for literary translation in low-resource settings—especially for genres demanding nuanced handling of narrative style and cultural meaning. Existing open and proprietary translation models are typically not tuned for creative content, and English–Romanian parallel datasets remain scarce, limited in scope, or non-public. Prior benchmarks overwhelmingly favor news or technical content, leaving literary and didactic narratives comparatively neglected.
@@
Building on the English-only fable generation efforts of DS-TF1-EN-3M [18], we systematically extend to cross-lingual translation, focusing on moral fables—short stories that require not only semantic accuracy but also fluency, narrative coherence, stylistic faithfulness, and nuanced cultural adaptation. Generating high-quality parallel data under resource constraints is especially challenging for such genres, motivating our use of LLMs for both dataset creation and the development of an automated, multi-dimensional evaluation framework.
To address the challenges outlined above, our work makes the following key contributions:
  • Narrative-Aware Evaluation Protocol. We show how to bootstrap high-quality literary data with instruction-tuned LLMs, combining corpus-level BLEU [25] with a five-dimension LLM rubric for reproducible, fine-grained assessment.
@@
  • Three open, fine-tuned translation models. We release three open, fine-tuned translation models—tf2-1b, tf2-4b, and tf2-12b—LoRA-adapted checkpoints that substantially close the quality gap to proprietary systems while remaining efficient enough to run locally on commodity GPUs. [^3] https://huggingface.co/klusai/tf2-1b,  https://huggingface.co/klusai/tf2-4b,  https://huggingface.co/klusai/tf2-12b
  • Fully transparent artifacts. All data, code, prompts, and evaluation scripts are published under permissive licenses to foster reproducibility and further research in low-resource literary NLP.
By focusing on literary translation in a genuinely low-resource language, TF2 fills a gap left by existing datasets and MT models, demonstrating that high-quality, culturally adapted translation is feasible within tight cost and resource constraints.
Building on this foundation, the remainder of the paper is organized to systematically explore the practical and methodological questions that arise in the creation and evaluation of large-scale literary translation resources for low-resource languages. We motivate our experimental design, comparative analyses, and evaluation framework in the context of both real-world constraints and the broader landscape of open and proprietary language technologies.
@@ 1.1 Research Questions
In formulating our research questions, we recognize that the development of large-scale translation datasets and models is subject to a variety of practical constraints. Among these, financial cost remains the most immediate and stringent limitation, especially for low-resource language communities or research teams with restricted budgets. However, other important factors must also be considered, including the environmental impact associated with large-scale computation—such as CO2 emissions and water use for hardware cooling—as well as the ethical provenance of training data, particularly with respect to data privacy and fair representation. While our primary focus in this work is on optimizing inference for a domain-specific NLP task, we acknowledge that by doing so we also address sustainable and responsible machine translation research, hence helping to reduce ecological footprint and adhere to ethical data standards. These broader considerations motivate the structure of our experimental pipeline and inform our discussion of the trade-offs inherent in large-scale synthetic data generation.
Accordingly, we structure our study around the following central research questions (RQs):
@@ 1. RQ1
    Cost-Constrained Data Generation:
  2. RQ1
    Cost-Constrained Data Generation: Can a large-scale, high-quality English–Romanian literary translation dataset (moral fables) be built under strict budget constraints—orders of magnitude below state-of-the-art (SOTA) proprietary LLMs—using open models of 12B parameters or fewer, and what trade-offs are necessary to achieve this?
@@ 3. RQ2
    Open vs. Proprietary Translation Quality: How do SOTA proprietary models (such as GPT-o3 and Google’s Gemini) compare to open-source models (before and after fine-tuning) in terms of Romanian literary translation quality? Can a compact open model, fine-tuned on the generated fable corpus, approach the translation fidelity and fluency of much larger proprietary systems?
@@ 4. RQ3
    Evaluation Rigor and Automation: Can an automated evaluation framework using an LLM-as-a-judge reliably assess literary translation quality across multiple dimensions (accuracy, fluency, coherence, style, cultural adaptation)? How well do these LLM-based evaluations align with expectations of professional literary translation, and what are the limitations of this approach?
By addressing RQ1, we aim to establish a practical pipeline for low-cost dataset creation. RQ2 examines the effectiveness of budget-constrained open models relative to the best available closed models, shedding light on the trade-offs between accessibility and performance. RQ3 focuses on ensuring that our evaluation methodology is robust and can capture the nuances of literary translation better than crude metrics like BLEU alone.
@@ 2 Related Work
Large language models have been widely used to fabricate training corpora for tasks where human data are scarce or costly. Early work on back-translation [22] showed that automatically generated target–source pairs can boost neural MT in low-resource regimes. More recently, LLM-driven pipelines have produced massive monolingual or parallel datasets at a fraction of traditional annotation cost [14, 31]. In narrative domains, Eldan and Li [4] created TinyStories, demonstrating that millions of synthetic tales enable coherent story generation with sub-10M-parameter models. Our TF2 corpus follows this line of research but focuses on bilingual literary content—an area where synthetic data remain under-explored.
@@
Moral or didactic stories have attracted interest as test-beds for value-aligned generation and commonsense reasoning. Guan et al. [6] introduced STORAL, a human-authored dataset coupling narratives with morals; subsequent work explored moral inference and story completion. TinyFabulist DS-TF1-EN-3M [18] scaled synthetic moral fables to three million English examples. TF2 extends this effort cross-lingually, supplying the first large-scale English–Romanian parallel fable corpus and highlighting translation-specific challenges such as cultural adaptation of morals.
@@
Comparative studies of closed APIs (e.g., GPT-4, Gemini) and open-weight models (Llama, Mistral, Qwen) report a measurable but narrowing quality gap, especially for less-represented languages [36]. Parameter-efficient fine-tuning methods—most notably LoRA adapters [7] and subsequent PEFT toolkits [32]—have demonstrated substantial gains for open models, closing much of the gap on task-specific benchmarks while maintaining low deployment costs. Recent open-weight multilingual giants such as DeepSeek-LLM [1], EuroLLM-9B [16], and Meta’s NLLB-200 [27] achieve impressive translation quality across dozens to hundreds of languages, including Romanian. However, their large parameter counts and memory requirements place them out of reach for many budget- or hardware-constrained users. By contrast, TF2 focuses on parameter-efficient adaptation of compact and mid-sized open models (1B, 4B, and 12B parameters), demonstrating that with careful data curation, instruction tuning, and adapter compression, even relatively lightweight models can approach the translation quality of much larger proprietary systems for creative literary tasks. This three-pronged strategy supports scalable, cost-effective deployment, while still yielding strong empirical results in both automated and rubric-based evaluation.
@@
While BLEU [20] remains the de-facto MT benchmark, its correlation with human judgment drops for creative text. Frameworks such as MQM [13] and COMET [21] address adequacy and fluency more directly, but require either expert annotators or supervised reference models. Recent work leverages LLMs as automatic judges, achieving human-level agreement for summarisation and dialogue evaluation [12]. We adopt this LLM-as-judge paradigm, extending it with a five-dimension rubric tailored to literary translation—accuracy, fluency, coherence, style, and cultural adaptation.
Several recent approaches push evaluation further in ways highly relevant to narrative translation. Hu et al. [8] propose Cont-COMET, a context-aware extension of COMET that incorporates preceding and following sentences to better align with human annotations, improving both system-level and segment-level assessments. This context-sensitive design highlights how broader narrative coherence could also be integrated into evaluation—a direction we aim to extend in future work. Wang et al. [29] introduce a multi-agent system where different LLMs assess terminology consistency, narrative perspective, and stylistic fidelity, combining their judgments into an overall translation quality score. Zhang et al. [35] present LiTransProQA, a QA-based evaluation framework where LLMs answer targeted questions about style, tone, and cultural references, mapping responses into quantitative scores. Finally, Yao et al. [34] release the CAMT corpus alongside metrics focused on culture-specific items, showing that LLM-based evaluation outperforms neural MTs in cultural adaptation.
@@
Finally, a growing body of work emphasises budget-constrained model development, from efficient inference engines [5] to dataset generation strategies that trade minimal API spend for maximal downstream benefit [30]. TF2 contributes an end-to-end case study: dataset synthesis, fine-tuning, and evaluation—all within a moderate budget by current standards—thus offering a practical blueprint for researchers with limited compute or funding.
@@
Recent efforts have started to address the scarcity of open benchmarks and models for Romanian natural language processing. Masala et al. [17] present OpenLLM-Ro, a comprehensive suite of open-source Romanian language models, evaluation benchmarks, and datasets, released through both GitHub [^4] https://github.com/openllm-ro and Hugging Face [^5] https://huggingface.co/OpenLLM-Ro. These resources provide a foundation for research in Romanian LLMs, but, to our knowledge, do not specifically target literary translation or creative narrative domains. Our TF2 benchmark is complementary, focusing on literary translation tasks and on the cross-lingual transfer of narrative content, thereby addressing a different—yet crucial—aspect of Romanian language technology.
Taken together, these strands of research underscore the feasibility and importance of cost-efficient, open-model pipelines for literary translation in low-resource languages. Our work contributes to this growing field by introducing TF2 based on an approach that combines synthetic data generation, parameter-efficient fine-tuning, and narrative-aware evaluation to produce high-quality English–Romanian literary translations under strict budget constraints.
@@ 3 TinyFabulist Translation Framework (TF2)
TF2 generalizes the original TinyFabulist Framework (TF) project to a bilingual and scalable setting, introducing an end-to-end methodology for constructing a large-scale, high-quality English–Romanian literary translation resource. The framework is architected for transparency, replicability, and practical cost-awareness at every step, emphasizing open data, parameter-efficient adaptation, and rigorous evaluation.
The workflow (see Figure 1) consists of the following four main stages, each described and justified in detail below:
@@
    Evaluating Candidate Translators: We benchmark 13 diverse LLMs and commercial translation APIs on a held-out fable set (e.g. in our case the DS-TF1-EN-3M dataset [18]), with evaluation targeting not only accuracy but also literary style and coherence. Metrics include a five-dimension LLM-based rubric, capturing the specific requirements of creative translation.
@@
    Parallel Dataset Creation: The best-performing Stage 1 system is used to automatically translate 15.000 English fables, producing the TinyFabulist DS-TF2-EN-RO-15K corpus. This “silver-standard” parallel set supplies essential supervision for downstream model adaptation in a domain lacking human references, and also introduces a ground-truth reference enabling BLEU-based benchmarking.
@@
    Parameter-Efficient Fine-Tuning: A suite of open LLMs—ranging from 1B to 12B parameters—are fine-tuned on the 15k parallel set using Low-Rank Adaptation (LoRA [7]), yielding domain-specialized English–Romanian translation models. Multiple quantized variants are produced for efficient local inference. The fine-tuning process is designed to be transparent and reproducible.
@@
    Large-Scale Corpus Generation: The best fine-tuned model(s) are used to translate the remaining $\sim$3M English fables from DS-TF1-EN-3M dataset, resulting in the largest openly available bilingual literary dataset for Romanian. The pipeline is modular, enabling ongoing evaluation and future expansion as models or translation strategies improve.
Each stage is tightly integrated with rigorous quality control and open resource release, providing a replicable blueprint for low-resource literary MT beyond English–Romanian.
@@
Figure 1: TinyFabulist Translation Framework (TF2) pipeline: Evaluation of translation models on literary benchmarks (S1); creation of a 15k English–Romanian parallel corpus via best system (S2); parameter-efficient fine-tuning of open LLMs and quantized variants (S3); and large-scale Romanian translation of the full English corpus (S4).

### v2

# Building Large-Scale English–Romanian Literary Translation Resources with Open Models

@@ 1 Introduction
Translating literature into low-resource languages poses unique challenges, given both the scarcity of parallel data and the need to faithfully preserve stylistic nuance and cultural context. Romanian—a Latin language spoken by over 24 million people—remains notably underserved in the realm of literary translation resources. Existing machine translation (MT) benchmarks, such as WMT [^1] WMT stands for the Conference on Machine Translation (formerly “Workshop on Machine Translation”). The WMT organizes annual shared tasks for benchmarking machine translation systems, with its “news translation” task being the most widely used standard evaluation for translation of contemporary news articles. See: https://www.statmt.org/wmt23/., predominantly focus on informational or news text, thus providing limited guidance for creative, narrative-driven translation tasks. Recent advances in large language models (LLMs) have enabled synthetic text generation at unprecedented scale, opening new opportunities for augmenting MT datasets [14, 19]. Notably, projects such as TinyStories [4] demonstrate that even compact models can generate coherent narratives when trained on carefully curated synthetic corpora. However, much of the existing work has focused on monolingual story generation in English [6, 4] or on improvements to translation via back-translation [22], rather than on the systematic creation of large-scale bilingual literary datasets. As a result, open questions remain regarding how both proprietary and open-source LLMs can be leveraged to construct and evaluate literary translations in genuinely low-resource settings.
@@
Substantial progress has also been made in building large bilingual and multilingual resources. For example, CLIRMatrix [26] compiles parallel data across 139 languages for cross-lingual retrieval, and mT5 [33] pretrains a text-to-text transformer on a wide range of language pairs and domains. These resources, however, primarily target general-domain or retrieval-oriented MT and include limited literary content. Even for language pairs such as English–Romanian, coverage is typically restricted to news, web, or miscellaneous domains. To date, no openly available, large-scale English–Romanian parallel corpus has been specifically curated for creative narratives or fables.
@@
TinyFabulist Translation Framework (TF2) addresses this gap by extending our earlier TinyFabulist-TF1 corpus of English-only fables [18] into a bilingual setting. The resulting dataset, DS-TF2-EN-RO-3M, comprises three million synthetic English–Romanian parallel fables, accompanied by fine-tuned open-source translation models and a literary translation benchmark. By combining parameter-efficient model adaptation with a narrative-oriented evaluation protocol, TF2 provides both a large-scale resource and a practical experimental framework for studying literary machine translation in a low-resource language pair.
@@
Recent multilingual models such as NLLB-200 [27] and EuroLLM [16] achieve strong performance across many language pairs, but are generally optimized for breadth and general-domain translation. Their large parameter counts can also limit deployment in cost-constrained or on-device scenarios. In our experiments, these models do not consistently capture the stylistic nuance, discourse coherence, and cultural adaptation required for literary translation (Section 4), motivating the exploration of more targeted, domain-adaptive approaches.
@@
Building on the English-only fable generation efforts of DS-TF1-EN-3M [18], we focus on moral fables—short narratives that require semantic fidelity alongside fluency, narrative coherence, stylistic faithfulness, and culturally appropriate moral framing. Generating parallel data for such content under resource constraints motivates the use of LLMs for both dataset construction and scalable evaluation.
Our contributions are summarized as follows:
  • Narrative-Aware Evaluation Protocol. We introduce a multi-dimensional evaluation framework for literary translation that combines corpus-level BLEU [25] with an LLM-based rubric assessing accuracy, fluency, coherence, style, and cultural adaptation. In this work, BLEU is used strictly as a secondary consistency check, while primary analysis relies on rubric-based and human-anchored evaluation.
@@
  • Three open, fine-tuned translation models. We release three LoRA-adapted English–Romanian translation models—tf2-1b, tf2-4b, and tf2-12b—which substantially narrow the performance gap to proprietary systems under cost-aware evaluation, while remaining efficient enough to run on commodity GPUs. [^3] https://huggingface.co/klusai/tf2-1b,  https://huggingface.co/klusai/tf2-4b,  https://huggingface.co/klusai/tf2-12b
  • Fully transparent artifacts. All datasets, code, prompts, and evaluation scripts are released under permissive licenses to support reproducibility and further research in low-resource literary NLP.
@@ 1.1 Research Questions
The construction of large-scale translation datasets and models is constrained by multiple practical considerations. Financial cost is often the most immediate limitation, particularly for low-resource language communities or research teams with restricted budgets. Additional factors include the environmental impact of large-scale computation—such as CO2 emissions and water use for hardware cooling—and the ethical provenance of training data, especially with respect to data privacy and fair representation. These considerations motivate the design of our experimental pipeline and inform the trade-offs examined in this work.
Accordingly, we structure our study around the following research questions:
@@ 1. RQ1
    Cost-Constrained Data Generation: Can a large-scale English–Romanian literary translation dataset be constructed under strict budget constraints using small open models, and what trade-offs does this entail?
@@ 2. RQ2
    Open vs. Proprietary Translation Quality: How do fine-tuned open models compare to proprietary systems for English–Romanian literary translation under cost-aware and resource-constrained settings?
@@ 3. RQ3
    Evaluation Rigor and Automation: To what extent can an automated, LLM-based evaluation framework provide useful and interpretable signals for literary translation quality, when anchored by expert human judgment on a representative subset?
The remainder of this paper is organized as follows. Section 2 reviews related work on literary machine translation, synthetic data generation, and evaluation. Section 3 describes the TF2 pipeline in detail. Section 4 presents experimental results and comparative evaluations. Section 5 introduces the released datasets and metadata schema. Sections 6 and 7 discuss findings and limitations, respectively.
@@ 2 Related Work
Large language models have been widely used to generate training corpora for tasks where human-annotated data are scarce or costly. Early work on back-translation [22] demonstrated that automatically generated target–source pairs can substantially improve neural MT in low-resource settings. More recent LLM-driven pipelines have enabled the creation of massive monolingual and parallel datasets at a fraction of traditional annotation cost [14, 31]. In narrative domains, Eldan and Li [4] introduced TinyStories, showing that large collections of synthetic narratives can support coherent story generation even with compact models. Despite these advances, synthetic data for bilingual literary translation remain comparatively underexplored.
@@
Moral and didactic narratives have attracted interest as test-beds for studying value-aligned generation and commonsense reasoning. Guan et al. [6] introduced STORAL, a human-authored dataset pairing narratives with explicit moral lessons, which has since been used for tasks such as moral inference and story understanding. Subsequent work has explored scaling narrative datasets and using synthetic generation to model moral reasoning in text. These efforts highlight the suitability of short narrative forms, such as fables, for studying higher-level discourse phenomena, including coherence, moral framing, and cultural interpretation.
@@
Comparative studies of proprietary APIs (e.g., GPT-4, Gemini) and open-weight models (e.g., LLaMA, Mistral, Qwen) report a persistent but narrowing quality gap, particularly for less-represented languages and domains [36]. Parameter-efficient fine-tuning methods, most notably LoRA adapters [7] and subsequent PEFT toolkits [32], have been shown to substantially improve task-specific performance of open models while keeping training and deployment costs low. Recent multilingual open-weight models such as DeepSeek-LLM [1], EuroLLM-9B [16], and Meta’s NLLB-200 [27] demonstrate strong general-purpose translation performance across many languages, including Romanian. However, their scale and hardware requirements can limit accessibility in cost- or resource-constrained settings, motivating continued interest in compact, adaptable translation models.
@@
BLEU [20] remains a widely used reference-based metric for MT evaluation, but its correlation with human judgment is known to decrease for creative and stylistically rich text. Alternative frameworks such as MQM [13] and COMET [21] provide more fine-grained assessments of adequacy and fluency, though they typically require expert annotations or supervised quality estimation models. More recently, LLMs have been employed as automatic judges, showing promising agreement with human evaluations in tasks such as summarization and dialogue assessment [12].
Several recent approaches further extend evaluation in directions relevant to narrative translation. Hu et al. [8] propose Cont-COMET, a context-aware extension of COMET that incorporates surrounding sentences to better capture discourse-level effects. Wang et al. [29] introduce a multi-agent evaluation framework in which multiple LLMs assess terminology consistency, narrative perspective, and stylistic fidelity. Zhang et al. [35] present LiTransProQA, a question-answering-based evaluation method targeting style, tone, and cultural references. Finally, Yao et al. [34] release the CAMT corpus alongside metrics designed to evaluate culture-specific translation phenomena, demonstrating the importance of culturally aware evaluation signals.
@@
A growing body of work emphasizes budget-constrained model development and deployment, ranging from efficient inference and quantization techniques [5] to dataset generation strategies that trade limited API usage for downstream performance gains [30]. These efforts illustrate how careful design choices can enable competitive NLP systems under practical cost and hardware constraints.
@@
Recent initiatives have begun to address the scarcity of open benchmarks and models for Romanian NLP. Masala et al. [17] introduce OpenLLM-Ro, a collection of open-source Romanian language models, datasets, and evaluation benchmarks released via GitHub [^4] https://github.com/openllm-ro and Hugging Face [^5] https://huggingface.co/OpenLLM-Ro. While these resources provide a valuable foundation for Romanian language technology, they primarily target general-purpose language understanding and generation rather than literary translation or creative narrative domains.
Taken together, prior work highlights the potential of synthetic data, parameter-efficient adaptation, and cost-aware evaluation strategies for advancing machine translation in low-resource settings. These insights motivate the pipeline and resources introduced in the following sections.
@@ 3 TinyFabulist Translation Framework (TF2)
TF2 generalizes the original TinyFabulist Framework (TF) project to a bilingual and scalable setting, introducing an end-to-end methodology for constructing an English–Romanian literary translation resource. The framework emphasizes transparency, replicability, and cost-awareness, prioritizing open artifacts, parameter-efficient adaptation, and evaluation that combines automatic metrics, rubric-based assessment, and targeted human judgment.
The workflow (see Figure 1) consists of four stages:
@@
    Evaluating Candidate Translators: We benchmark 13 diverse LLMs and commercial translation APIs on a held-out fable set drawn from DS-TF1-EN-3M [18]. Evaluation targets multiple aspects of literary translation and uses a five-dimension LLM-based rubric to compare systems in a scalable and consistent manner.
@@
    Parallel Dataset Creation: The top-ranked Stage 1 system is used to translate 15,000 English fables, producing the TinyFabulist DS-TF2-EN-RO-15K corpus. This silver-standard parallel dataset provides supervision for downstream adaptation in a setting without human-authored references, and supplies a consistent reference corpus for subsequent reference-based evaluation (e.g., BLEU).
@@
    Parameter-Efficient Fine-Tuning: A suite of open LLMs—ranging from 1B to 12B parameters—are fine-tuned on the 15k parallel set using Low-Rank Adaptation (LoRA; 7), yielding domain-adapted English–Romanian translation models. Quantized variants are produced to support efficient local inference.
@@
    Large-Scale Corpus Generation: The best-performing fine-tuned model(s) are used to translate the remaining $\sim$3M English fables from DS-TF1-EN-3M, resulting in a large bilingual English–Romanian corpus for Romanian literary MT. The pipeline is modular and supports future re-generation as translation models or evaluation protocols improve.
Each stage is integrated with quality control and open artifact release, providing a reproducible blueprint for low-resource literary MT beyond English–Romanian.
@@
Figure 1: TinyFabulist Translation Framework (TF2) pipeline: evaluation of translation models on literary benchmarks (S1); creation of a 15k English–Romanian parallel corpus via the top-ranked system (S2); parameter-efficient fine-tuning of open LLMs and quantized variants (S3); and large-scale translation of the full English corpus (S4).